OpenStudy (anonymous):

Choose the replacement for x and y that makes this equation a true statement: 2x + 55y = −214

OpenStudy (anonymous):

sounds like a multiple choice question?

OpenStudy (anonymous):

There are many many many correct answers, but in a multiple choice question, there is probably only one ...

OpenStudy (anonymous):

yes here are the choices : A: x=-3 , y = -4 B: x=-4 , y= 3 C: x=3 , y=-4 D: x=4 y: 4

OpenStudy (anonymous):

Try each one. I will do the first one for you Substitute x = -3 and y = -4 into 2x + 55y = -214. The work is shown below: 2(-3) + 55(-4) = -214 or -6 - 220 = -214 -226 = -214 Do these match? .......... no ...... the answer is not A

OpenStudy (anonymous):

Did you try b?

OpenStudy (anonymous):

the answer is d because b was 157 and c was -215 .

OpenStudy (anonymous):

almost ..... d?? Substitute x = 4 and y = 4 into 2x + 55y = -214. The work is shown below: 2(4) + 55(4) = -214 or 8 + 220 = -214 or 228 = -214 Do these match? .......... no ...... the answer is not D

OpenStudy (anonymous):

B or C?